land Find the derivate of each function write it in simplified form without negative exponents. b. f(x) = 4x sinx²
Answer #1

we have

f(x)=4x^3sin(x^2)

f'(x)=\frac{d}{dx}4x^3sin(x^2)

f'(x)=4x^3\frac{d}{dx}sin(x^2)+sin(x^2)\frac{d}{dx}4x^3

f'(x)=4x^3cos(x^2)\frac{d}{dx}(x^2)+sin(x^2)(12x^2)

f'(x)=4x^3cos(x^2)(2x)+12x^2sin(x^2)

f'(x)=8x^4cos(x^2)+12x^2sin(x^2)